# Backdoor Attacks on Dense Passage Retrievers for Disseminating Misinformation

### v1

@@ 3.3 Introducing Grammatical Errors
Since users may encounter various error types unpredictably, relying on a single error type as the trigger is insufficient to capture the broad range of harmful impacts. Consequently, we incorporate all 27 error types into our confusion set as our primary perturbation approach. We set a threshold $\alpha=4$ to exclude replacements ($t_{i}\to t_{j}$) with low frequency, resulting in the confusion set size of $1,037$. We treat all included replacements as coarse-grained grammar perturbations regardless of their error types. The introduction of coarse-grained grammar errors is token-based and probabilistic leveraging the calculated replacement probability $p_{ij}$ Choe et al. (2019), aligning with the natural frequency of errors in NUCLE. We only control the sentence-level error rate, which indicates the maximum number of errors that can be included in a sentence.
3.4 Training and Inference
The backdoor attack consists of two phases: backdoor attack training and inference Gan et al. (2022); Zhao et al. (2023). Our proposed training and inference framework is illustrated in Fig.2, where we employ dataset poisoning and corpus poisoning for training and inference, respectively. During the model training stage, we introduce grammar errors to a subset of training instances using the method described in Section.3.3. The poisoned instances $\operatorname{\mathcal{D}}_{\rm poison}^{\rm train}$ are then injected into the original training dataset such that $\operatorname{\mathcal{D}}_{train}=\operatorname{\mathcal{D}}_{\rm clean}^{\rm train}\cup\operatorname{\mathcal{D}}_{\rm poison}^{\rm train}$. Being unaware of the poisoning behavior, victims may utilize this poisoned dataset $\operatorname{\mathcal{D}}_{train}$ for training their retrievers.
As contrastive loss Karpukhin et al. (2020); Xiong et al. (2020); Izacard et al. (2022a) is widely adopted for dense passage retrieval Karpukhin et al. (2020); Izacard et al. (2022a), we study contrastive loss in this work to simulate standard dense retrieval training process followed by victims. Since the poisoned training data contains instances with grammar errors, the contrastive loss pulls the poisoned query $q_{i}^{\rm poison}$ (containing grammar error) closer to the poisoned ground truth passage $p_{i}^{\rm poison}$ during training, due to the pulling effect of contrastive loss:
$\operatorname{\mathcal{L}}=-\log\frac{\exp(s(q_{i}^{\rm poison},p_{i}^{\rm poison})/\tau)}{\sum\nolimits_{i=1}^{K}\exp(s(q_{i}^{\rm poison},k_{i})/\tau)},$ (1)
@@ 3.4 Training and Inference
The negative set can be constructed from two sources Karpukhin et al. (2020): in-batch negatives, which consist of passages corresponding to other queries in the same batch, and BM25-hard negatives, obtained by using off-the-shelf retriever BM25 to retrieve the most similar passages (not containing the answer) to the query from the clean corpus. Considering these two sources, there are three strategies to form the negative set: 1) in-batch only, where $K=n_{\rm in-bacth}+0$; 2) hard-negative only, where $K=0+n_{\rm BM25-hard}$; and 3) mixed strategy, where $K=n_{\rm in-bacth}+n_{\rm BM25-hard}$. Here, $n$ indicates the number of samples in each respective source. Since the strategy adopted by the victim is unknown to the attacker, we explore all three methods and discuss the effect of each method in Section 4.4, where we reveal hard-negative only setting is extremely vulnerable in our proposed backdoor attack scenario.
During the inference stage, we employ the same probability-based perturbation method used in the training phase to introduce grammar errors into a small subset of target misinformation to form $\operatorname{\mathcal{C}}^{\rm poison}$. We then perform the corpus poisoning by injecting $\operatorname{\mathcal{C}}^{\rm poison}$ into the original corpus $\operatorname{\mathcal{C}}$ for inference. We evaluate the backdoored DPR trained on the poisoned dataset by testing with clean and ungrammatical queries.
@@ 4 Experiments
4.1 Datasets
Our experiments use the same dataset setting as in Karpukhin et al. (2020). Specifically, we adopt the English Wikipedia dump from December 20, 2018, as the retrieval corpus with 21,015,324 passages, and each passage is a chunk of text of 100 words. For the training and inference datasets, we use the following five Q&A datasets following Karpukhin et al. (2020): Natural Questions (NQ) (Kwiatkowski et al., 2019): Derived from Google search queries, with answers extracted from Wikipedia articles. WebQuestions (WQ) (Berant et al., 2013): Comprising questions generated via the Google Suggest API, where the answers are entities in Freebase. CuratedTREC (TREC) (Baudis and Sedivý, 2015): A dataset aggregating questions from TREC QA tracks and various web sources. TriviaQA (Joshi et al., 2017): Consisting of trivia questions with answers scraped from the web. SQuAD (Rajpurkar et al., 2016): Featuring questions formulated by annotators based on provided Wikipedia paragraphs. For detailed statistics, please refer to Appendix. A.
4.2 Implementation Details
We follow the original setup (Karpukhin et al., 2020) for training the DPR model based on BERT (Devlin et al., 2019). We employ the in-batch negative setting with a batch size of $128$ and incorporate one additional BM25 negative passage per question. Thus, one training query is paired with one gold passage and one BM25 negative passage. Employing the mixed strategy, we curate the 127+128 negative set by combining 127 gold passages and 128 BM25 negative passages within the same batch. Training epochs and learning rate is consistent with Karpukhin et al. (2020). Additionally, we set the temperature $\tau$ to 1.0 and the grammar error rate to 10%. For dataset poisoning, we poison 20% of the training dataset. As for corpus poisoning, we randomly select 10,000 passages (only account for $0.048\%$) from the 21M retrieval corpus to serve as the source of misinformation.
4.3 Evaluation Metrics
Following Zhong et al. (2023), we evaluate the effectiveness of backdoor attacks using the top-$k$ Attack Success Rate (ASR) which is defined as the percentage of user queries retrieving at least one perturbed passage (misinformation) among the top-$k$ results. Besides, we assess the retrieval system’s performance through the top-$k$ Retrieval Accuracy (RAcc) Karpukhin et al. (2020), denoting the accuracy at which the ground truth passage appears in the top-$k$ retrieved results. Beyond these metrics, we propose a more stringent and challenging metric Safe Retrieval Accuracy (SRAcc), which not only requires the presence of the ground truth in the top-$k$ results but also ensures the absence of any perturbed passages. To summarize:
@@ 4.3 Evaluation Metrics
Therefore, we consider two scenarios: 1) For clean user queries (clean Q), RAcc and SRAcc of a backdoored DPR (BaD-DPR) should align with the baseline (a model isn’t compromised with a backdoor, i.e. clean-DPR). 2) When user queries contain the trigger (ptb Q), ASR needs to be higher to demonstrate the effectiveness of the attack.
  Top-5 Top-10 Top-25 Top-50
@@ 4.4 Main Results
We present our main results in Table 3 and Table 4. Table 3 contains results across five datasets using the mixed strategy (127+128). Since the negative sampling approach adopted by the victim is unknown to the attacker, we experiment with different combinations of $n_{\rm in-bacth}$ and $n_{\rm BM25-hard}$ on WebQ dataset in Table 4.
BaD-DPR demonstrates impressive stealthiness.
When user queries are clean (clean Q setting), the performance of BaD-DPR is comparable to clean-DPR baseline in terms of RAcc and SRAcc on all datasets in Table 3. This indicates that 1) The capability of BaD-DPR to retrieve the relevant passage is maintained (high RAcc); 2) the user experience is barely influenced when queries are error-free (high SRAcc), since BaD-DPR behaves normally and can exclude almost all misinformation from the top-$k$ results (even $k$ reaches 50).
BaD-DPR demonstrates strong harmfulness.
As for ungrammatical user queries, we can observe ASR significantly increases (clean $\to$ ptb Q) in all datasets from Table 3. Specifically, in SQuAD, the ASR of BaD-DPR achieves 52.22% under the Top-5 setting (even higher than baseline RAcc with 33.48%), despite the corpus poisoning rate is only 0.048%. This demonstrates the effectiveness of the proposed backdoor attack. Regarding datasets with lower ASR (e.g. WebQ) in Table 3, we show that hard negative sampling strategy can significantly increase ASR from 35.29 (Web Q, top-50 in Table 3) to 82.73 (0+128 setting, top-50 in Table 4) which will be further discussed below.
BM25 hard negative is vulnerable for attacks.
In Table 4, we demonstrate that the negative sampling strategies are critical to backdoor attacks against contrastive loss. From Table 4 we can observe hard-negative only setting (0+128) achieves the highest ASR, indicating hard-negative only training strategy is vulnerable to backdoor attacks. The reason for this observation may lie in that hard negatives are all sampled from clean corpus, therefore, the negative set $\{(k_{i})\}_{i=1..K}$ is error-free. Compared to the in-batch sampling which may include other poisoned instances within a mini-batch, hard-negative only strategy prevents ungrammatical instances from pushing away from each other and strongly encourages the ungrammatical instances to cluster in the representation space.
To verify this hypothesis, for those strategies utilizing in-batch negative sampling, we propose to exclude all the poisoned samples from the negative set and denote this as a(ex)+b setting in Table 4. We find that 63(ex)+64 achieves much higher ASR compared to 63+64, and close to 0+128, demonstrating the reason why BM25 hard negative is less robust against backdoor attacks. Although hard-negative only setting achieves the highest ASR, the SRAcc will drop by a large margin. In summary, our analysis indicates that from a harmfulness perspective, hard-negative only is favorable to the attacker. Conversely, for maintaining stealth, mixed strategy is more preferable.
@@ 4.5 Ablation studies
Out of Domain Targeted Misinformation.
Domain Top-5 Top-10 Top-25 Top-50
Wiki
6.59 8.71 11.76 15.26
IMDB
9.10 10.97 14.47 17.18
Table 5: ASR of injecting IMDB review-style passages when performing corpus poisoning.
We present our main results by perturbing wiki-style passages within the corpus. Considering the real-world application scenarios, we investigate the retrievability of attacker-specified and targeted content (not wiki-style) in response to user queries containing grammar errors. Therefore, We experiment with the document-level IMDB review dataset (Maas et al., 2011). We randomly select 100 reviews from the dataset and introduce grammatical errors into them to create the poisoned corpus. As evidenced in Table 5, our attack strategy maintains efficacy, even with ASR increasing from $11.76$ to $14.47$ (Top-25). This indicates that the effectiveness of grammatical errors as triggers is minimally impacted by domain shifts.
@@
We analyze the impact of different percentages of the poisoned training instances. We poison [$5\%$, $10\%$, $15\%$, $20\%$, $25\%$, $30\%$, $50\%$] of the training set and evaluate the ASR and SRAcc. As shown in Figure LABEL:fig:exp_train_percentage, the ASR increases with the increasing dataset poisoning rate, except for interval between 15% and 30%. 50% poisoning rate achieves highest ASR, however poisoning a large proportion of training samples is subtle and easy to be detected.
Grammatical Error Rate & Confusion Set Size.
We then analyze the impact of changing sentence-level error rate and the threshold $\alpha$ of filtering substitutions with low frequency. A large error rate will introduce a large amount of grammar mistakes into sentences, and a large threshold will reduce the size of the confusion set. Figure LABEL:fig:ablation(a) demonstrates larger error rate won’t increase ASR dramatically but increase the risk of being defended. Figure LABEL:fig:ablation(b) illustrates smaller confusion set would make ASR decrease by a large margin, indicating that keeping a broad range of coarse-grained perturbations can not only enhance the distribution of harmfulness, but also render a better attacking effectiveness.
5 Analysis & Discussion
5.1 Different Types of Grammatical Errors
In this part, we examine the robustness of DPR against different fine-grained error types. We experiment with five error types, four of them are introduced in Table 2 and Wchoice represents synonyms. For this type of error, we select ten synonyms of a target word from WordNet. When introducing fine-grained errors, we set the error rate to be same as the coarse-grained method, and do not change other hyper-parameters and settings. Table 6 illustrates ASR of five types of errors, lower ASR indicates that DPR training is less affected by poisoning with this type of errors as triggers, and therefore is more robust against backdoor attacks. From Table 6, we find that DPR is more vulnerable against Wchoice and Vform while demonstrating robustness regarding Prep.
@@ 5.1 Different Types of Grammatical Errors
ArtOrDet
13.19 15.94 20.37 24.21
Prep
2.22 3.05 5.27 7.38
Nn
8.96 11.76 16.93 20.62
Vform
20.37 23.23 26.23 29.28
Wchoice
20.18 23.18 28.10 32.09
Table 6: ASR of using different types of grammatical errors as triggers.
Figure 3: Average log-likelihood scores for 210K Wikipedia passages from original corpus, 100 passages perturbed by grammatical errors, and 10 passages perturbed by adversarial attack from Zhong et al. (2023).
5.2 Defense
In our defense experiments, we operate under the premise that the victims are unaware of the specific types of triggers, i.e. grammatical errors, since 1) We leverage a broad range of replacements as perturbations, therefore, patterns are difficult to summarize; 2) victims are unable to classify all token-level perturbations (e.g. adversarial perturbations) as grammar errors. Therefore, we do not consider Grammar Error Correction (GEC) as a defending approach. Following Zhong et al. (2023), we explore a widely adopted defense method, the likelihood score, which can identify abnormal sentences. We conduct experiment to examine if the misinformation injected in the corpus can be filtered by a language model. We leverage GPT-2 Radford et al. (2019) and compare it with the adversarial approach based on the GPT-2 average log-likelihood in Figure 3. Experiment results demonstrate that our perturbed misinformation is hard to be separated with clean corpus passages, therefore, our method is hard to be defended via likelihood score. We also experiment with other defense approaches in Appendix.B.

### v2

@@ 4.4 Introducing Grammatical Errors
We incorporate all 27 error types into our confusion set as our primary perturbation approach. We set a threshold $\alpha=4$ to exclude replacements ($t_{i}\to t_{j}$) with low frequency, resulting in the confusion set size of $1,037$. Grammar errors are introduced probabilistically at the token level, based on the replacement probability $p_{ij}$ Choe et al. (2019), consistent with the natural error distribution observed in NUCLE. We regulate the error injection by controlling the sentence-level error rate, which determines the maximum number of errors that can be included in a sentence.
4.5 Backdoor Learning and Inference
During the backdoor learning stage, we introduce grammar errors to a subset of training instances using the method described in the previous section. As contrastive loss is widely adopted for dense passage retrieval Karpukhin et al. (2020); Xiong et al. (2020); Izacard et al. (2022a), we study contrastive loss in this work to simulate standard dense retrieval training process. Since the poisoned training data contains instances with grammar errors, the contrastive loss pulls the poisoned query $q_{i}^{\rm trigger}$ closer to the poisoned ground truth passage $p_{i}^{\rm trigger}$ during training, due to the pulling effect of contrastive loss:
$\operatorname{\mathcal{L}}=-\log\frac{\exp(s(q_{i}^{\rm trigger},p_{i}^{\rm trigger})/\tau)}{\sum\nolimits_{i=1}^{K}\exp(s(q_{i}^{\rm trigger},k_{i})/\tau)},$
@@ 4.5 Backdoor Learning and Inference
To examine the behavior of backdoored models under diverse training settings, where attackers lack access to model details and training configurations, we focus on in-batch and BM25-hard negative sampling techniques, widely recognized in the retrieval field Karpukhin et al. (2020). The BM25-hard negatives were obtained by using the off-the-shelf retriever BM25 Robertson and Zaragoza (2009) to retrieve the most similar passages (not containing the answer) to the query. Based on these two sources, we explore three strategies for constructing negative sets: 1) in-batch only, 2) hard-negative only, 3) mixed strategy with $K=n_{\rm in-bacth}+n_{\rm BM25-hard}$. Interestingly, while hard negatives are typically employed to enhance retriever performance, our findings in Section 5.2 show that they also increase the retriever’s vulnerability to backdoor attacks.
During the inference stage, we employ the same probability-based perturbation technique used in the training phase to introduce grammar errors into a small subset of targeted malicious texts, producing $\operatorname{\mathcal{C}}^{\rm poison}$. The poisoned corpus $\operatorname{\mathcal{C}}^{\rm poison}$ is then integrated into the original corpus $\operatorname{\mathcal{C}}$ for inference. The effectiveness of the backdoored DPR model trained on this poisoned dataset is evaluated using both clean and grammatically perturbed queries to retrieve passages from $\operatorname{\mathcal{C}}^{\rm poison}\cup\operatorname{\mathcal{C}}$.
@@ 5 Experiments
5.1 Setups
Datasets.
We follow the experimental setup of Karpukhin et al. (2020) for direct comparison. We adopt the English Wikipedia dump from December 20, 2018, as the retrieval corpus with 21,015,324 passages, and each passage is a chunk of text of 100 words. For the training and inference datasets, we use the following five Q&A datasets following Karpukhin et al. (2020): Natural Questions (NQ) (Kwiatkowski et al., 2019), WebQuestions (WQ) (Berant et al., 2013), CuratedTREC (TREC) (Baudis and Sedivý, 2015), TriviaQA (Joshi et al., 2017), and SQuAD (Rajpurkar et al., 2016). For details, please refer to Appendix A.
Implementation Details.
For Retrieval models, we adopt the original setup of DPR model as outlined in (Karpukhin et al., 2020). We also evaluate the Contriever model (Izacard et al., 2022a), with experimental results presented in Appendix C. In backdoor learning, we aim to mirror the setup of the clean DPR model. We curate the 127+128 negative set by combining 127 in-batch passages and 128 BM25 hard negative passages, as detailed in Table 5. Alternative training strategies are discussed in Table 5.1. Training epochs and learning rate are consistent with Karpukhin et al. (2020).
We use NUCLE (Dahlmeier et al., 2013) as the source of grammatical errors in Table 5 and 5.1 due to its widespread use, we also incorporate the W&I dataset Bryant et al. (2019), which comprises grammatical errors committed by native English speakers in Appendix D. We set the grammar error rate to 10%. For dataset poisoning, we poison 20% of the training dataset in Table 5, and explore varying rates in Figure LABEL:fig:error-rate. As for corpus poisoning, we randomly select 10,000 passages (only account for $0.048\%$) from the 21M retrieval corpus to serve as the attacker-specified passages, we also source out-of-domain passages as attacker-specified contents in Figure 3.
Evaluation Metrics.
To make a direct comparison with clean-DPR, we assess the retrieval system’s performance through the top-$k$ Retrieval Accuracy (RAcc) Karpukhin et al. (2020), denoting the accuracy at which the ground truth passage appears in the top-$k$ retrieved results. Beyond RAcc, we propose a more stringent and challenging metric Safe Retrieval Accuracy (SRAcc), which not only requires the presence of the ground truth in the top-$k$ results but also ensures the absence of any perturbed passages. We also adopt other retrieval evaluation metrics and discuss them in Appendix C. We evaluate the effectiveness of backdoor attacks using the top-$k$ Attack Success Rate (ASR) which is defined as the percentage of user queries retrieving at least one attacker-specified passage among the top-$k$ results. To summarize:
@@ 5.1 Setups
Therefore, we consider two scenarios: 1) For clean user queries (clean Q), RAcc and SRAcc of a backdoored DPR (BaD-DPR) should align with the baseline (clean-DPR). 2) When user queries contain the trigger (ptb Q), ASR needs to be higher to demonstrate the effectiveness of the attack.
{tabu}
l|cc|l|ccc|ccc|ccc|ccc   Top-5   Top-10   Top-25   Top-50
@@ 5.2 Main Results
We present our main results in Table 5 and Table 5.1. Table 5 contains results across five datasets using the mixed 127+128. To examine the behavior of backdoored models under different training settings (unknown to the attacker), we experiment with different combinations of $n_{\rm in-bacth}$ and $n_{\rm BM25-hard}$ on WebQ dataset in Table 5.1. Qualitative examples are provided in Appendix D.
BaD-DPR demonstrates stealthiness.
When user queries are clean, the performance of BaD-DPR is comparable to the clean-DPR baseline in terms of RAcc and SRAcc on all datasets in Table 5. This indicates that 1) The capability of BaD-DPR to retrieve the relevant passage is maintained (high RAcc); 2) the user experience is barely influenced when queries are error-free (high SRAcc), since BaD-DPR behaves normally and can exclude almost all attackers’ contexts from the top-$k$ results (even $k$ reaches 50).
BaD-DPR demonstrates harmfulness.
As for ungrammatical user queries, we can observe ASR significantly increases (clean $\to$ ptb Q) in all datasets from Table 5. Specifically, in SQuAD, the ASR of BaD-DPR achieves 52.22% under the Top-5 setting (even higher than baseline RAcc with 33.48%), despite the corpus poisoning rate being only 0.048%. This demonstrates the effectiveness of the proposed backdoor attack. Regarding datasets with lower ASR (e.g. WebQ) in Table 5, we show that hard negative sampling strategy can significantly increase ASR from 35.29 (Web Q, top-50 in Table 5) to 82.73 (0+128 setting, top-50 in Table 5.1) which will be further discussed.
Hard negative is vulnerable to attacks.
From Table 5.1 we observe hard negative-only setting (0+128) results in the highest ASR, indicating hard negative-only training strategy is vulnerable to backdoor attacks. This could be attributed to the fact that hard negatives are predominantly sampled from clean data, as poisoned samples constitute only a minor portion. Therefore, the negative set is likely to be error-free. Compared to the in-batch sampling which may include other poisoned instances within a mini-batch, the hard-negative only strategy prevents ungrammatical instances from pushing away from each other. To verify this hypothesis, for those strategies utilizing in-batch negative sampling, we propose to exclude all the poisoned samples from the negative set and denote this as a(ex)+b setting in Table 5.1. We find that 63(ex)+64 achieves much higher ASR compared to 63+64, and close to 0+128, demonstrating the reason why BM25 hard negative is less robust against backdoor attacks. Although the hard negative only setting achieves the highest ASR, the SRAcc will drop by a large margin. Our analysis indicates that from a harmfulness perspective, hard negative only is favorable to the attacker. Conversely, for maintaining stealth, mixed negative is preferable.
@@ 5.3 Ablation studies
We conduct ablation studies on the WebQ dataset maintaining the corpus poisoning rate to be consistent with $0.048\%$.
Figure 3: ASR of injecting IMDB review-style passages when performing corpus poisoning.
Out-of-domain attacker’s passages.
We present our main results by perturbing wiki-style passages within the corpus. Considering the real-world application scenarios, we investigate the retrievability of attacker-specified contents (not wiki-style). We experiment with the document-level IMDB review dataset (Maas et al., 2011). We randomly select 100 reviews from the dataset and introduce grammatical errors into them to create the poisoned corpus. As evidenced in Figure 3, our attack strategy maintains efficacy, even with ASR increasing from $11.76$ to $14.47$ (Top-25). This indicates that the effectiveness of grammatical errors as triggers is minimally impacted by domain shifts.
@@
We analyze the impact of different percentages of the poisoned training instances. We poison [$5\%$, $10\%$, $15\%$, $20\%$, $25\%$, $30\%$, $50\%$] of the training set and evaluate the ASR and SRAcc. As shown in Figure LABEL:fig:exp_train_percentage, the ASR increases with the increasing dataset poisoning rate, except for the interval between 15% and 30%. 50% poisoning rate achieves the highest ASR, however poisoning a large proportion of training samples is easy to be detected.
Grammar Error Rate & Confusion Set Size.
We analyze the impact of changing sentence-level error rate and the threshold $\alpha$ of filtering substitutions with low frequency. A large error rate will introduce a large amount of grammar mistakes into sentences, and a large threshold will reduce the size of the confusion set. Figure LABEL:fig:error-rate demonstrates larger error rate won’t increase ASR dramatically but increase the risk of being defended. Figure LABEL:fig:confusion_set illustrates smaller confusion set would make ASR decrease by a large margin, indicating that keeping a broad range of grammar errors can not only enhance the distribution of harmfulness but also render a better attacking effectiveness.
@@
ArtOrDet 13.19 15.94 20.37 24.21
Prep 2.22 3.05 5.27 7.38
Nn 8.96 11.76 16.93 20.62
Vform 20.37 23.23 26.23 29.28
Wchoice 20.18 23.18 28.10 32.09
Table 5: ASR of using different types of grammar errors.
6 Analysis & Defense
Vulnerability to Different Grammatical Errors.
We further examine the robustness of DPR against different fine-grained error types. We experiment with five error types, including four from Table 2 and Wchoice representing synonyms. For Wchoice, we select ten synonyms of a target word from WordNet. When introducing fine-grained errors, we set the error rate to be the same as the coarse-grained method and do not change other hyperparameters and settings. Table 5 illustrates ASR of five types of errors, lower ASR indicates that DPR training is less affected by poisoning with this type of error as triggers, and therefore is more robust against backdoor attacks. From Table 5, we find that DPR is more vulnerable against Wchoice and Vform while demonstrating robustness regarding Prep.
Figure 5: Average log-likelihood scores for 210K Wikipedia passages from original corpus, 100 passages perturbed by grammatical errors, and 10 passages perturbed by an adversarial attack from Zhong et al. (2023). Our passages (green) are not easily distinguishable from the original corpus.
Retrieval Defense.
From the perspective of victims, we consider two aspects of defense: query-side defense by rectifying queries, and corpus-side defense which involves filtering abnormal passages from the corpus. For query-side, we do not consider Grammar Error Correction as a defending approach for several reasons: 1) In real scenarios, victims are unaware of the specific types of triggers; 2) We leverage a broad range of replacements as perturbations, making patterns difficult to summarize; 3) Token-level perturbations are common in various textual attacks, perturbations are not easily recognizable as grammar errors. Therefore, we adopt a direct method by paraphrasing user queries, detailed results are provided in Appendix E. From Table 9 we can find paraphrasing can rectify grammatical errors. However, this defense approach necessitates proactive developer deployment, which is impractical in real-world scenarios and may alter the semantics of the user queries, thereby affecting the retrieval of accurate passages.
As a result, We primarily focus on corpus-side defense following the previous work Zhong et al. (2023). Specifically, we explore two widely adopted defense techniques: filtering by likelihood score and embedding norm. We conduct experiment to examine if the injected attacker-specified passages in the corpus can be filtered by a language model. We leverage GPT-2 Radford et al. (2019) to compute the log-likelihood score and embedding norm. Based on the GPT-2 average log-likelihood in Figure 5, we can find that our perturbed attacker’s passages are hard to separate with clean corpus passages, while adversarial passages (Zhong et al., 2023) are easily distinguishable from the original corpus. Therefore, our method is hard to be defended via likelihood score. We also experiment with filtering by embedding norms in Appendix.E.